### v1

# Robust Predictive Motion Planning by Learning Obstacle Uncertainty

Abstract
Safe motion planning for robotic systems in dynamic environments is nontrivial in the presence of uncertain obstacles, where estimation of obstacle uncertainties is crucial in predicting future motions of dynamic obstacles. The worst-case characterization gives a conservative uncertainty prediction and may result in infeasible motion planning for the ego robotic system. In this paper, an efficient, robust, and safe motion-planing algorithm is developed by learning the obstacle uncertainties online. More specifically, the unknown yet intended control set of obstacles is efficiently computed by solving a linear programming problem. The learned control set is used to compute forward reachable sets of obstacles that are less conservative than the worst-case prediction. Based on the forward prediction, a robust model predictive controller is designed to compute a safe reference trajectory for the ego robotic system that remains outside the reachable sets of obstacles over the prediction horizon. The method is applied to a car-like mobile robot in both simulations and hardware experiments to demonstrate its effectiveness.
Index Terms:
Robust motion planning, predictive control, uncertainty quantification, safe autonomy.
I Introduction
Figure 1: Reach-avoid scenario where the ego vehicle (EV) aims at planning a safe trajectory in the presence of an uncertain surrounding vehicle (SV). (a) Deterministic approach: no consideration of motion uncertainty of the obstacle, therefore the planning is unsafe. (b) Worst-case robust approach: considering the worst-case motion uncertainty of the obstacle yields an infeasible problem. (c) The proposed approach: reducing the conservatism and planning a feasible reference trajectory.
Safe motion planning is an important research topic in control and robotics [1], [2]. Dynamic and uncertain environments make it challenging to perform efficient, safe and robust planning for a robotic system. One crucial source of uncertainty from the environment is obstacle uncertainty, which refers to the uncertain future motion of dynamic obstacles [3]. This uncertainty can be exemplified by unknown decision-making of human-driven vehicles, which typically acts as uncertainties to autonomous vehicles [4].
Predicting obstacle uncertainties plays a key role in predicting future motions of dynamic obstacles [5]. Although the worst-case realization of uncertainties in theory ensures safety of the ego robotic system, it results in overly conservative solutions or even infeasibility of the motion-planning problem [4], [6], [7]. To reduce the conservativeness from the worst-case uncertainty characterization, this paper develops a safe motion-planning algorithm that learns the obstacle uncertainties using real-time observations of the obstacles.
The motivation can be highlighted by the reach-avoid scenario in Fig. 1, where an autonomous ego vehicle (EV) navigates to reach the target while avoiding a surrounding vehicle (SV), e.g., a human-driven vehicle, in the driveable area. It is assumed that there is no direct communication between the two vehicles. Since the intended trajectory of the SV is unknown to the EV, the key point is how to predict the motion of the SV. A simple way for the prediction is by propagating a constant velocity model. Despite the computational advantage, the predicted trajectory of the SV may in this way have a deviation from its intended trajectory, as shown in Fig. 1(a). Thus, such predictions could be insufficient for the EV to perform safe motion planning. Alternatively, assuming that the EV knows the SV’s dynamics and (possibly overestimated) admissible control set, it is reasonable to predict the maximal forward occupancy that contains all trajectory realizations of the SV [^1] Forward occupancy represents the set of positions a system can reach over a specific horizon, defining a region that could be occupied based on the system’s control capabilities.. Although uncertainties of the SV are incorporated in the prediction, the overly conservative forward occupancy may make the motion-planning problem infeasible, as illustrated in Fig. 1(b). Therefore, the primary objective of this paper is to refine the robust prediction method to reduce its conservativeness while maintaining safety. This involves understanding the intended control set of the SV, which, though typically unknown, can be learned from data sampled by observations of SV. By computing the forward occupancy of the SV based on learning its intended control set, the EV can plan a trajectory that is both safe and feasible, as shown in Fig. 1(c).
To realize the objective, this paper considers the motion planning of a general ego robotic system in the presence of dynamic surrounding obstacles with control strategies unknown to the ego system. A new efficient way is proposed to learn the decision-making support, i.e., the set of intended control actions of each obstacle, using observations of obstacles. Thus, the use of this set enables a better motion prediction, in the sense of less conservative than the worst-case prediction. The main contributions of this paper are:
  • (a)
    A novel approach is proposed to learn the unknown yet intended control set of obstacles without making any assumptions about the distribution of control actions of the obstacles and without the need for training with prior data. The set is efficiently computed by solving a linear programming (LP) problem.
  • (b)
    Using the online learned set, a robust predictive motion planner is developed for motion planning of the ego system subject to collision avoidance with uncertain surrounding obstacles in dynamic environments. The method can perform resilient motion planning in the absence of prior knowledge regarding obstacle uncertainties.
  • (c)
    The performance of the proposed motion-planning method is validated through both simulations and hardware experiments in several benchmark traffic scenarios.
Outline
This paper is organized as follows: Section II places this research in the context of related work in the literature. Section III formalizes the general research problem. Section IV specifies the method for learning the control set of obstacles. Section V designs the robust MPC for safe motion planning based on uncertainty prediction. Sections VI and VII show the performance of the method in both simulations and hardware experiments, and Section VIII concludes the paper.
II Related Work
This section reviews the most related research from two perspectives: (1) Uncertainty prediction of dynamic systems, and (2) uncertainty-aware motion planning in time-varying environments.
II-A Uncertainty Prediction
The robust approach for predicting uncertainties of a dynamic system is by formulating a sequence of sets over the horizon to cover the worst-case uncertainty realizations. This is typically achieved by forward reachability analysis [8]. For example, [2] characterized the forward reachable sets (FRSs) of the ego robotic system considering the worst-case disturbances for safety guarantees in robust motion planning. In [9], the FRSs were predicted for both an autonomous ego vehicle and a surrounding vehicle based on the maximal disturbance set for safety verification. Focusing on uncertainty predictions of other traffic participants, the later work in [10] and [11] combined the worst-case FRSs with the road networks and the interactions with the autonomous ego vehicle, respectively, in practical driving scenarios. An alternative yet more conservative approach to the forward reachability analysis is computing the minimal robust positive invariant (MRPI) set for the ego system under the maximal disturbances, and relevant works can be found in [12], [13], and [14].
A main problem of robust methods is that the worst-case assumptions can lead to unnecessary conservatism in the prediction. To mitigate this problem, the non-conservative methods, which formulate the set of uncertainties as a subset of the worst-case set, have been developed. For instance, [15] and [16] formulated the probabilistic reachable set of other vehicles in traffic by assuming the decision-making process of surrounding vehicles, which is influenced by the road geometry, as a Markov chain. If sufficient prior knowledge of obstacle uncertainties is available, e.g., a large amount of training data, the probabilistic prediction can be performed by machine learning models like neural ordinary differential equations [17]. The training data can also be utilized to formulate the empirical reachable set that contains the $\alpha$-likely observed trajectories of a system [18]. In addition, with particular assumptions on the distribution of obstacles, a confidence-aware prediction that formulated an obstacle occupancy with a degree of prediction confidence/risk was designed in [19] and [4], where the former assumed that the control action of an obstacle satisfies a Boltzmann distribution, and the later assumed that the leading vehicle in the automated overtaking scenario respects a supermartingale. Apart from the probabilistic prediction, learning the true disturbance set of a dynamic system was also investigated to reduce the conservativeness, as presented in [20] that learned the disturbance set of an ego helicopter using residuals, and [21] that learned the disturbance set of an ego vehicle using a rigid tube.
II-B Uncertain-Aware Motion Planning
Based on the prediction results, the motion-planning problems in uncertain and dynamic environments can be solved using various well-established methods [22]. Among them, model predictive control (MPC) stands out because of the straightforward inclusion of differential constraints of the system and safety constraints with uncertain obstacles and the environment [23]. Combined with different uncertainty-prediction methods, the MPC-based planners include robust MPC, stochastic MPC, and scenario MPC [24]. For example, in [3] and [25], the robust MPC was used to generate the safe reference trajectory for an autonomous vehicle subject to the worst-case obstacle occupancy. In contrast, the stochastic MPC, as presented in [26] and [27], considers collision avoidance with obstacle occupancy with a probability based on the assumptions on the uncertainty distribution of the obstacle. The scenario MPC can also achieve similar probabilistic constraints, where the constraints are transformed into deterministic forms using finite samples of uncertainties [28]. In addition to these methods, there have been combinations of them, like the robust scenario MPC that optimizes policies concerning multi-mode obstacle-uncertainty predictions [29], and the scenario stochastic MPC to also handle multi-modal obstacle uncertainties with known distributions [30].
III Problem Statement
This section first defines the general notations used in the paper, then introduces models of the ego system and dynamic surrounding obstacles and finally formulates the robust motion-planning problem to be solved.
III-A Notations
$\mathbb{R}^{n}$ is the $n$-dimensional real number vector space, $\mathbb{R}_{+}^{n}$ is the $n$-dimensional non-negative vector space, $\mathbb{N}^{n}$ is the $n$-dimensional natural number vector space, and $\mathbb{N}_{+}^{n}$ is the $n$-dimensional positive integer vector space. $[0,1]^{n}$ means an $n$-dimensional vector with each element between $[0,1]$. $I^{n}$ indicates an $n\times n$ identity matrix. Matrices of appropriate dimension with all elements equal to 1 and 0 are denoted by $\bm{1}$ and $\bm{0}$, respectively. An interval of integers is denoted by $\mathbb{I}_{a}^{b}=\{a,a+1,\cdots,b\}$. For two sets $\mathcal{A}$ and $\mathcal{B}$, $\mathcal{A}\oplus\mathcal{B}=\left\{x+y\ |\ x\in\mathcal{A},y\in\mathcal{B}\right\}$. For a set $\Gamma\subset\mathbb{R}^{n_{a}+n_{b}}$, the set projection is defined as $\mathcal{A}={\rm Proj}_{a}(\Gamma)=\left\{a\in\mathbb{R}^{n_{a}}:\exists b\in\mathbb{R}^{n_{b}},(a,b)\in\Gamma\right\}$. For the MPC iteration, the current time step is indicated by $t$. A prediction of a variable $x$ at time step $t+i$ over the prediction horizon is represented as $x_{i|t}$, where $i=1,\ldots,N$, and $N$ is the prediction horizon. The sampling interval is represented as $T$. The cardinality of a set $\mathcal{A}$ is represented by $|\mathcal{A}|$.
III-B Modeling of Ego System and Surrounding Obstacles
The ego system’s dynamics are described by a discrete-time control system
$x^{e}_{t+1}=f(x^{e}_{t},\ u^{e}_{t}),$ (1)
where $x_{t}^{e}\in\mathbb{X}^{e}\subseteq\mathbb{R}^{n^{e}_{x}}$ is the state and $u_{t}^{e}\in\mathbb{U}^{e}\subseteq\mathbb{R}^{n^{e}_{u}}$ is the control input at time step $t$. Here, the superscript $e$ indicates the ego system. The sets $\mathbb{X}^{e}$ and $\mathbb{U}^{e}$ are the state set and admissible control set, respectively.
Consider an environment with multiple surrounding obstacles, where the motion of the center of geometry of each obstacle is modeled by a linear time-varying (LTV) system
$x^{s}_{t+1}=A^{s}_{t}x^{s}_{t}+B^{s}_{t}u^{s}_{t},$ (2)
where $A^{s}_{t}$ and $B^{s}_{t}$ are the system matrices at time step $t$, $x^{s}_{t}\in\mathbb{R}^{n^{s}_{x}}$ is the state and $u^{s}_{t}\in\mathbb{R}^{n^{s}_{u}}$ is the input at time step $t$. The superscript $s$ indicates the $s$-th obstacle and the index set of all involved obstacles is denoted by a finite set $\mathcal{S}$.
As illustrated in the motivating example in Section I, the control capability of the obstacles is distinguished by two different control sets. For the $s$-th obstacle, the admissible control set $\mathbb{U}^{s}$ refers to the set of control inputs determined by the physical limitations in the worst case, while the intended control set $\tilde{\mathbb{U}}_{t}^{s}\subseteq\mathbb{U}^{s}$ refers to the set that covers the control inputs that would be used by the obstacle from step $t$ over a planning horizon. It is assumed that $\mathbb{U}^{s}$ is known, or can be overly estimated, by the ego system, while $\tilde{\mathbb{U}}_{t}^{s}$ is unknown.
The sets $\tilde{\mathbb{U}}_{t}^{s}$ and $\mathbb{U}^{s}$ can be illustrated by the following example. The maximal acceleration of a vehicle at the limit of friction could be ±$\pm 8\ {\rm m/s^{2}}$. In practice, because of the capacity of the power system or caution of the driver, the vehicle may only generate accelerations within ±$\pm 2\ {\rm m/s^{2}}$. In this case, the admissible control set is $\mathbb{U}^{s}=[-8,8]\ {\rm m/s^{2}}$, while the intended control set is $\tilde{\mathbb{U}}_{t}^{s}=[-2,2]\ {\rm m/s^{2}}$, $\forall t\in\mathbb{N}$.
Example III.1 will present the models of the ego system and surrounding obstacles in a specific scenario to illustrate the general forms in (1) and (2).
Remark III.1.
The LTV model (2) is used for several reasons. First, a linear model can easily be derived to approximate the motion characteristics of the obstacle. Second, linear models for the obstacles facilitate efficient computation of forward reachable sets when predicting the motion of obstacles.
Example III.1.
Vehicle models in the reach-avoid problem: Consider the reach-avoid example in Fig. 1 throughout the paper, the EV in Fig. 1 is modeled by a nonlinear single-track kinematic model [27],
$\dot{p}_{x}^{e}$ $=v^{e}{\rm cos}\left(\varphi^{e}+\beta^{e}\right),$ (3a)
$\dot{p}_{y}^{e}$ $=v^{e}{\rm sin}\left(\varphi^{e}+\beta^{e}\right),$ (3b)
$\dot{\varphi}^{e}$ $=\frac{v^{e}}{l_{r}}{\rm sin}\left(\beta^{e}\right),$ (3c)
$\dot{v}^{e}$ $=a^{e},$ (3d)
$\dot{a}^{e}$ $=\eta^{e},$ (3e)
$\beta^{e}$ $={\rm arctan}\left(\frac{l_{r}}{l_{f}+l_{r}}{\rm tan}(\delta^{e})\right),$ (3f)
where $(p^{e}_{x},p^{e}_{y})$ is the coordinate of the center of geometry of the vehicle in the ground coordinate system. The variable $\varphi^{e}$ is the yaw angle, $\beta^{e}$ is the angle of the velocity of the center of geometry with the longitudinal axis of the vehicle, $v^{e}$, $a^{e}$, and $\eta^{e}$ mean the longitudinal speed, acceleration, and jerk, in the vehicle frame, respectively, $\delta^{e}$ is the front tire angle, and parameters $l_{f}$ and $l_{r}$ are the distances from the center of geometry to the front axle and rear axle, respectively. The model takes $a^{e}$ and $\eta^{e}$ as inputs and the other variables as states. This continuous-time model can be approximated to a discrete-time model in (1) using a sampling interval $T$ with different discretization methods [31].
The actual model of the SV in Fig. 1 is unknown. From the perspective of the EV, the motion of the center of geometry of the SV is modeled by a linear time-invariant model
$x_{t+1}^{s}=A^{s}x_{t}^{s}+B^{s}u_{t}^{s},$ (4)
where
$x_{t}^{s}=[p_{x,t}^{s}\ v_{x,t}^{s}\ p_{y,t}^{s}\ v_{y,t}^{s}]^{\rm T},\ u_{t}^{s}=[a_{x,t}^{s}\ a_{y,t}^{s}]^{\rm T},$
$A^{s}=\begin{bmatrix}1&T&0&0\\ 0&1&0&0\\ 0&0&1&T\\ 0&0&0&1\end{bmatrix},\ B^{s}=\begin{bmatrix}T^{2}/2&0\\ T&0\\ 0&T^{2}/2\\ 0&T\end{bmatrix},$
and the variables $p_{x,t}^{s}$, $v_{x,t}^{s}$, $a_{x,t}^{s}$ mean the longitudinal position, velocity, and acceleration at time step $t$, respectively, in the ground coordinate system. Variables $p_{y,t}^{s}$, $v_{y,t}^{s}$, and $a_{y,t}^{s}$ mean the lateral position, velocity, and acceleration at time step $t$, respectively, in the ground coordinate system. The control input satisfies $u_{t}^{s}\in\tilde{\mathbb{U}}_{t}^{s}\subseteq\mathbb{U}^{s}$, where the details will be given in case studies in Section VI.
III-C Research Problem
The goal of the ego system is to plan an optimal and safe reference trajectory to fulfill desired specifications in the presence of uncertain surrounding obstacles. The planned reference trajectory is anticipated to be less conservative than that of the worst-case realization, yet robust against the obstacles’ uncertain motions. Toward this goal, this paper has to solve three sub-problems: (i) learning the unknown intended control set $\tilde{\mathbb{U}}_{t}^{s}$ of each obstacle, i.e., computing a set to approximate $\tilde{\mathbb{U}}_{t}^{s}$; (ii) predicting the forward occupancy of each obstacle based on the obstacle model (2) and the learned intended control set of the obstacle; (iii) planning an optimal reference trajectory that enables the ego system to avoid collision with the occupancy of obstacles and satisfy differential and admissible constraints on the ego system model (1). The first two sub-problems are addressed in Section IV, and the motion-planning method is presented in Section V.
IV Learning Obstacle Uncertainty
This section describes how to learn the intended control set $\tilde{\mathbb{U}}_{t}^{s}$ of a general surrounding obstacle by a sampling-based approach, and how to predict the forward occupancy for the obstacle based on learning the set $\tilde{\mathbb{U}}_{t}^{s}$.
IV-A Learning Intended Control Set $\tilde{\mathbb{U}}_{t}^{s}$
Assumption IV.1.
At the current time step $t$, the ego system can measure or infer the control input $u^{s}_{t-1}$ of model (2) of the obstacle $s$, $\forall s\in\mathcal{S}$.
The estimation of the control input $u_{t-1}^{s}$ depends on observed information of the obstacle $s$ and the specific form of model (2). This will be specified in Remark VI.2 in the case studies. Since the intended control set is a hidden variable, an approach is to infer the hidden information using past observations [32]. To this end, based on Assumption IV.1, we define the information set $\mathcal{I}^{s}_{t}$ available to the ego system at time step $t$, which consists of the collected inputs of the $s$-th obstacle until time step $t$. The information set is updated as
$\mathcal{I}^{s}_{t}=\{u^{s}_{k}:k\in\mathbb{I}_{0}^{t-1}\}.$ (5)
The set $\mathcal{I}_{t}^{s}\subseteq\tilde{\mathbb{U}}_{t}^{s}$ contains observed control actions of obstacle $s$, and the process of generating the control actions by the obstacle is unknown to the ego system. Here, it is modeled as a random process such that the elements of the set $\mathcal{I}_{t}^{s}$ can be regarded as samples generated with unknown distributions from the intended control set of the obstacle. Next, the process of learning the unknown set $\tilde{\mathbb{U}}_{t}^{s}$ using the information set $\mathcal{I}^{s}_{t}$ and the known admissible control set $\mathbb{U}^{s}$ will be demonstrated. The following assumption is made before proceeding.
Assumption IV.2.
For any obstacle $s$, the admissible control set $\mathbb{U}^{s}$ is a convex and compact polytope that contains the origin in its interior, denoted by
$\mathbb{U}^{s}=\left\{u\in\mathbb{R}^{n^{s}_{u}}:H^{s}u\leq\bm{1}\right\},\ H^{s}\in\mathbb{R}^{n_{v}^{s}\times n_{u}^{s}}.$
In order to learn the unknown set $\tilde{\mathbb{U}}_{t}^{s}\subseteq\mathbb{U}^{s}$, a subset of $\mathbb{U}^{s}$ is parameterized in the form of
$\mathbb{U}^{s}(v,\rho,\bm{\theta})$ $=f(H^{s},\bm{\theta})\oplus(1-\rho)v,$ (6a)
$f(H^{s},\bm{\theta})$ =$=\left\{u\in\mathbb{R}^{n_{u}^{s}}:H^{s}u\leq\bm{\theta}\right\},$ (6b)
where $v\in\mathbb{U}^{s}$, $\rho\in[0,1]$, and $\bm{\theta}\in[0,1]^{n_{v}^{s}}$ are design parameters. It is shown in Lemma IV.1 that such parameterization under Assumption IV.2 constructs a subset of $\mathbb{U}^{s}$.
Lemma IV.1.
For any $v\in\mathbb{U}^{s}$, $\bm{\theta}\in[0,1]^{n_{v}^{s}}$, and $\rho\in[0,1]$, if $\bm{\theta}\leq\rho\bm{1}$, then $\mathbb{U}^{s}(v,\rho,\bm{\theta})\subseteq\mathbb{U}^{s}$.
Proof.
The explicit form of (6a)–(6b) can be written as
$\mathbb{U}^{s}(v,\rho,\bm{\theta})=\left\{u\in\mathbb{R}^{n^{s}_{u}}:H^{s}u\leq\bm{\theta}+(1-\rho)H^{s}v\right\}.$
It follows that $\bm{\theta}+(1-\rho)H^{s}v\leq\bm{\theta}+(1-\rho)\bm{1}\leq\bm{1}$. Therefore, $\mathbb{U}^{s}(v,\rho,\bm{\theta})\subseteq\mathbb{U}^{s}$. ∎
By parameterizing the set $\mathbb{U}^{s}(v,\rho,\bm{\theta})$ using (6a)–(6b) and leveraging its properties shown in Lemma IV.1, it is possible to find the optimal parameters $\rho^{\star}$, $v^{\star}$, and ${\bm{\theta}^{\star}}$ that minimize the volume of set $\mathbb{U}^{s}(v,\rho,\bm{\theta})$, while containing the historical information set $\mathcal{I}_{t}^{s}$. This optimal set, which is denoted by $\hat{\mathbb{U}}_{t}^{s}$, serves as the optimal approximation of the intended control set at time step $t$. The optimal parameters $v^{\star}$, $\rho^{\star}$, and ${\bm{\theta}^{\star}}$ can be obtained by solving the following problem:
$\mathop{\rm minimize}\limits_{v,\rho,\bm{\theta}}\quad$ $\bm{1}^{\rm T}\bm{\theta}+\rho$ (7)
${\rm subject\ to}\quad$ $H^{s}u^{s}-(1-\rho)H^{s}v\leq\bm{\theta},\ \forall u^{s}\in\mathcal{I}_{t}^{s},$
$H^{s}v\leq\bm{1},$
$\rho\in[0,1],\ \bm{\theta}\in[0,1]^{n_{v}^{s}},\ \bm{\theta}\leq\rho\bm{1},$
where the cost function minimizes the volume of set $\mathbb{U}^{s}(v,\rho,\bm{\theta})$ by minimizing summation of the vector ${\bm{\theta}}$ and its upper bound $\rho$. The first constraint of (7) implies that $\mathcal{I}_{t}^{s}\subseteq\mathbb{U}^{s}(v,\rho,\bm{\theta})$, i.e., the information set $\mathcal{I}_{t}^{s}$ is fully contained within the parameterized set $\mathbb{U}^{s}(v,\rho,\bm{\theta})$. The second and third constraints are required by Lemma IV.1. The problem (7) is non-convex as a result of the product of decision variables $v$ and $\rho$. The following proposition shows that (7) is equivalent to a linear programming (LP) problem.
Proposition IV.1.
Define an LP problem as
$\mathop{\rm minimize}\limits_{y,\rho,\bm{\theta}}\quad$ $\bm{1}^{\rm T}\bm{\theta}+\rho$ (8)
${\rm subject\ to}\quad$ $H^{s}u^{s}-H^{s}y\leq\bm{\theta},\ \forall u^{s}\in\mathcal{I}_{t}^{s},$
$H^{s}y\leq(1-\rho)\bm{1},$
$\rho\in[0,1],\ \bm{\theta}\in[0,1]^{n_{v}^{s}},\ \bm{\theta}\leq\rho\bm{1}.$
Denote the optimal solution of (8) by $(y^{\star},\rho^{\star},\bm{\theta}^{\star})$. Then, the optimal solution of (8) gives an optimal estimation of the unknown intended control set as
$\hat{\mathbb{U}}_{t}^{s}=f(H^{s},\bm{\theta}^{\star})\oplus y^{\star},$ (9)
where $f(H^{s},\bm{\theta}^{\star})$ is defined as in (6b). The set $\hat{\mathbb{U}}_{t}^{s}$ obtained in (9) is equivalent to the set obtained by solving (7).
Proof.
For (7) with $10\leq\rho<$ 1, it is equivalent to (8) with $10\leq\rho<$ 1. This equivalence arises from substituting the variable $y$ in (8) with $(1-\rho)v$. When $\rho=1$, (7) transforms into:
$\mathop{\rm minimize}\limits_{\rho,\bm{\theta}}\quad$ $\bm{1}^{\rm T}\bm{\theta}+\rho$ (10)
${\rm subject\ to}\quad$ $H^{s}u^{s}\leq\bm{\theta},\ \forall u^{s}\in\mathcal{I}_{t}^{s},$
$\rho=1,\ \bm{\theta}\in[0,1]^{n_{v}^{s}},$
where in (10), the constraint $H^{s}v\leq\bm{1}$ from (7) is removed as it is redundant. On the other hand, the equivalent form of (8) with $\rho=1$ is given by:
$\mathop{\rm minimize}\limits_{y,\rho,\bm{\theta}}\quad$ $\bm{1}^{\rm T}\bm{\theta}+\rho$ (11)
${\rm subject\ to}\quad$ $H^{s}u^{s}-H^{s}y\leq\bm{\theta},\ \forall u^{s}\in\mathcal{I}_{t}^{s},$
$H^{s}y\leq\bm{0},$
$\rho=1,\ \bm{\theta}\in[0,1]^{n_{v}^{s}}.$
Note that in (11), the constraint $H^{s}y\leq\bm{0}$ implies that $y=\bm{0}$. This confirms that (10) and (11) are equivalent with $\rho=1$. Hence, we conclude that (7) and (8) are equivalent, and the optimal quantified set $\hat{\mathbb{U}}_{t}^{s}$ is given in the form of (9). ∎
The sets $\mathbb{U}^{s}$, $\tilde{\mathbb{U}}_{t}^{s}$, and $\hat{\mathbb{U}}_{t}^{s}$ are illustrated in Fig. 2, which reflects that the set $\hat{\mathbb{U}}_{t}^{s}$ depends on the shape of $\mathbb{U}^{s}$ and the information set $\mathcal{I}_{t}^{s}$, and does not relate to the volume of $\mathbb{U}^{s}$.
Remark IV.1.
The formulation (8) is a standard randomized program [33], and it is shown in [33, Theorem 1] that the probability of intersection between $\hat{\mathbb{U}}_{t}^{s}$ and $\tilde{{\mathbb{U}}}_{t}^{s}$ is positively related to $|\mathcal{I}_{t}^{s}|$ if $\hat{\mathbb{U}}_{t}^{s}$ is invariant and the control action satisfies independent and identical distribution. While these assumptions are generally stringent in practice, the results highlight that the robustness of $\hat{\mathbb{U}}_{t}^{s}$ is enhanced with a larger information set $\mathcal{I}_{t}^{s}$.
Figure 2: Learning the set $\tilde{\mathbb{U}}_{t}^{s}$ by solving LP (8).
IV-B Computing $\hat{\mathbb{U}}_{t}^{s}$ with Low Complexity
Note that the computational complexity of the LP (8) for computing $\hat{\mathbb{U}}_{t}^{s}$ is polynomial in $|\mathcal{I}_{t}^{s}|$ if $\mathcal{I}_{t}^{s}$ is increasing with time. This means the complexity of (8) increases with each time step. To address this problem, two ways are proposed to manage the online computational complexity.
IV-B1 Moving-Horizon Approach
The first way is to control the size of $\mathcal{I}_{t}^{s}$ in a moving-horizon manner. That is, given a horizon $L\in\mathbb{N}_{+}$, the set $\mathcal{I}_{t}^{s}$ is updated as
$\mathcal{I}^{s}_{t}=\{u^{s}_{k}:k\in\mathbb{I}_{t-L-1}^{t-1}\}$
with an initialization $|\mathcal{I}^{s}_{0}|=L$. In this way, (8) is accompanied by $L$ scenario constraints.
IV-B2 Online Recursion
A second way is to perform an online recursive computation of $\hat{\mathbb{U}}_{t}^{s}$ based on $\hat{\mathbb{U}}_{t-1}^{s}$ and $u^{s}_{t-1}$. This is achieved by solving the following problem
$\mathop{\rm minimize}\limits_{v,\rho,\bm{\theta}}\quad$ $\bm{1}^{\rm T}\bm{\theta}+\rho$ (12)
${\rm subject\ to}\quad$ $u_{t-1}^{s}\in{\mathbb{U}}^{s}(v,\rho,\bm{\theta}),$
$\hat{\mathbb{U}}^{s}_{t-1}\subseteq\mathbb{U}^{s}(v,\rho,\bm{\theta}),$
$v\in\mathbb{U}^{s},$
$\rho\in[0,1],\ \bm{\theta}\in[0,1]^{n_{v}^{s}},\ \bm{\theta}\leq\rho\bm{1}.$
The optimal solution $(v^{\star},\rho^{\star},\bm{\theta}^{\star})$ of (12) enables update of $\hat{\mathbb{U}}_{t}^{s}$, which is the minimum set in the form of (6a) that covers the previous obtained set $\hat{\mathbb{U}}^{s}_{t-1}$ and the latest information sample $u_{t-1}^{s}$. Following the proof of Proposition IV.1, with $y=v(1-\rho)$, the optimal solution of the following LP provides the optimal estimation of the set $\tilde{\mathbb{U}}_{t}^{s}$ in the form of (9):
$\mathop{\rm minimize}\limits_{y,\rho,\bm{\theta}}\quad$ $\bm{1}^{\rm T}\bm{\theta}+\rho$ (13)
${\rm subject\ to}\quad$ $H^{s}u_{t-1}^{s}-H^{s}y\leq\bm{\theta},$
$H^{s}y^{\star}_{\rm pre}+\bm{\theta}^{\star}_{\rm pre}\leq H^{s}y+\bm{\theta},$
$H^{s}y\leq(1-\rho)\bm{1},$
$\rho\in[0,1],\ \bm{\theta}\in[0,1]^{n_{v}^{s}},\ \bm{\theta}\leq\rho\bm{1},$
where $y^{\star}_{\rm pre}$ and $\bm{\theta}^{\star}_{\rm pre}$ are the optimal solution of (13) at time step $t-1$, i.e., $\hat{\mathbb{U}}^{s}_{t-1}=f(H^{s},\bm{\theta}_{\rm pre}^{\star})\oplus y_{\rm pre}^{\star}$. The formulation (13) is computationally efficient as the number of constraints does not change with $|\mathcal{I}_{t}^{s}|$, and it does not lose any information when new samples are observed and added in the set $\mathcal{I}_{t}^{s}$.
IV-C Forward Reachability Analysis Based on $\hat{\mathbb{U}}_{t}^{s}$
Given $\hat{\mathbb{U}}_{t}^{s}$ obtained by (13), and the system described as in (2), forward reachability can be used to predict the occupancy of obstacles. Denote by $\hat{\mathcal{O}}_{i|t}^{s}$ the occupancy of the center of geometry of obstacle $s$ predicted $i$ steps ahead of the current time step $t$. The set $\hat{\mathcal{O}}_{i|t}^{s}$ can be computed as[4]
$\hat{\mathcal{O}}_{i|t}^{s}$ $={\rm Proj}_{\rm position}(\hat{\mathcal{R}}_{i|t}^{s}),$ (14a)
^$\hat{\mathcal{R}}_{i+1|t}^{s}$ $=A_{t+i}^{s}\hat{\mathcal{R}}^{s}_{i|t}\oplus B_{t+i}^{s}\hat{\mathbb{U}}_{t}^{s},$ (14b)
^$\hat{\mathcal{R}}_{0|t}^{s}$ =$=\left\{x_{t}^{s}\right\},$ (14c)
where ^$\hat{\mathcal{R}}_{i|t}^{s}$ is the forward reachable set at time step $t+i$ based on the set $\hat{\mathbb{U}}_{t}^{s}$, $x_{t}^{s}$ is the measured state of the obstacle at time step $t$, ${\rm Proj}_{\rm position}(\hat{\mathcal{R}}_{i|t}^{s})$ denotes the set of positions projected from the reachable set.
Figure 3: The learned intended control sets by formulation (8), formulation (13), the moving-horizon (MH) method, and the rigid tube (RT) method (for the MH method $|\mathcal{I}_{t}^{s}|=30$ ). (a) The number of samples is $50$. (b). The number of samples is $500$.
TABLE I: Computation Time of Different Methods for Learning the Set
Methods Computation time
Number of samples = $50$ Number of samples = $500$
Formulation (13) $2.5\ {\rm ms}$ $2.6\ {\rm ms}$
Formulation (8) $61.34\ {\rm ms}$ $695.8\ {\rm ms}$
Moving-horizon method $30.94\ {\rm ms}$ $28.9\ {\rm ms}$
Rigid tube method $1.6\ {\rm ms}$ $1.2\ {\rm ms}$
IV-D Performance Evaluation
This subsection evaluates the proposed method for learning the intended control set of a system, and the accuracy of the resultant forward reachability analysis.
IV-D1 Comparison Between Different Methods for Learning the Intended Control Set
The performance of formulation (13) for learning the intended control set of the obstacle is compared with the original formulation (8), the moving-horizon (MH) method that fixes $|\mathcal{I}_{t}^{s}|=L$ (see Section IV-B1), and the rigid tube (RT) method [21]. Suppose that the intended control set of a dynamic system is learned by observing the longitudinal and lateral accelerations in the ground coordinate system. The system takes the control action from a bounded but unknown set $\tilde{\mathbb{U}}_{t}^{s}$ with unknown distribution, and its admissible control set $\mathbb{U}_{t}^{s}$ is overly estimated as a regular hexagon. The results are shown in Fig. 3, and the computation time is summarized in Table I. The analyses through Fig. 3 and Table I reflect that the result by (13) closely aligns with the original formulation (8), yet requires less computation time. Compared to the RT method that tends to overestimate certain areas, the learned set by (13) is less conservative. It also surpasses the accuracy of the MH method, which tends to underestimate the set. Notably, the learned set expands when the number of samples increases, while the computation time remains in the same order. This implies an enhanced robustness of the result without compromising computational performance.
IV-D2 Analysis of the Responsiveness of the Method
This subsection analyzes the responsiveness of (13) in learning the control set of a system as in Section IV-D1. The system is designed to perform mild actions before time step $t=50$, and from time step $t=51$ it shifts to executing more aggressive maneuvers from an unknown set $\tilde{\mathbb{U}}_{t}^{s}$. The volume of the learned set with the discrete time steps, as well as the 2-norm of the sampled control action ($||u_{t}^{s}||$), are presented in Fig. 4. In addition, the learned control sets at different time steps are shown in Fig. 5. It is seen in Fig. 4 that after the system starts to take aggressive control actions from $t=51$, and the learned control set can rapidly respond to the change, the delay between the set volume and the change of the control style is just $3$ steps. Fig. 5 shows how the learned set starts from a minor initial set and grows when more samples are obtained. The set expands notably from $t=50$ to $t=54$ as a few aggressive control actions are sampled. The set at $t=100$ contains both the mild and aggressive control actions. This reflects that the method does not lose any information during the online learning process.
Figure 4: Control actions of the system and volume of learned control set.
Figure 5: Learned control sets at different time steps (at $t=50$ the mild control behavior ends, and at $t=100$ the sampling ends).
Figure 6: Predicted occupancy of a nonlinear vehicle model.
IV-D3 Reachability Analysis
This section evaluates the results of reachability analysis through in (IV-C) based on the learned control set $\hat{\mathbb{U}}_{t}^{s}$. The system in this case is modeled by a nonlinear kinematic single-track vehicle model as in (III.1), where the control inputs generated from an unknown input set $\tilde{\mathbb{U}}_{t}^{s}$ with unknown distribution. The accelerations in the global coordinate system are extracted through historical trajectories to learn its intended control set by solving (13). Then it uses (IV-C) to predict the occupancy over a horizon, where the dynamics of the system are modeled by a second-order integrator as in (4). It is seen in Fig. 6 that only a few samples are outside of the predicted occupancy, and the occupancy successfully covers the sampled trajectories at the end of the horizon. This indicates enhanced robustness over time. In addition, even though the nonlinear model of the system and the linear second-order integrator model used for predicting the occupancy in (IV-C) are quite different, the LP-based method still offers the desired accuracy in predicting the occupancy.
In addition, compared with a similar method, our proposed method has advantages in calculation time. Among the related work in Section II-A, the reachability analysis based on solving mixed-integer linear programming (MILP) problem [18] exhibits the most similar characteristics to the proposed method as it does not require the distribution of behaviors and model information of the system. The MILP-based method can get similar occupancy as in Fig. 6 by tuning the confidence level [18, Fig. 4]. For the same prediction horizon, the computation time by the proposed method with $100$ trajectories and $500$ trajectories is $0.0296\ {\rm s}$ and $0.0304\ {\rm s}$. For the MILP-based method, as presented in [18, Table I], the average computation time with $100$ trajectories and $500$ trajectories is $0.0322\ {\rm s}$ and $0.1352\ {\rm s}$. This reflects that our proposed method has a better computational performance than the MILP-based method, and the computation time remains unaffected by the number of samples.
V Robust Motion-Planning Strategy
This section formulates the robust MPC for safe motion planning of the ego system, based on the predicted obstacle occupancy $\hat{\mathcal{O}}_{i|t}^{s}$ defined in (14a). During the online motion-planning process, at each time step $t$, the ego system updates the set $\hat{\mathbb{U}}_{t}^{s}$ and the predicted occupancy $\hat{\mathcal{O}}_{i|t}^{s}$, and then solves the following optimal control problem (OCP)
$\mathop{\rm minimize}\limits_{u^{e}_{i-1|t}}\quad$ ${V}\left(x^{e}_{i|t},\ u^{e}_{i-1|t},\ {r}\right)$ (15a)
${\rm subject\ to}\quad$ $x^{e}_{i|t}=f(x^{e}_{i-1|t},u^{e}_{i-1|t}),$ (15b)
$x^{e}_{i|t}\in\mathbb{X}^{e},$ (15c)
$u_{i-1|t}^{e}\in\mathbb{U}^{e},$ (15d)
${\rm dis}(p^{e}_{i|t},\hat{\mathcal{O}}_{i|t}^{s})>d^{s}_{\min},\ \forall s\in\mathcal{S},$ (15e)
where $i=1,\ldots,N$. In the cost function (15a), ${r}$ denotes the reference state. The cost function ${V}(\cdot)$ is defined for measuring the performance of the motion planning. The constraints (15b)–(15d) concerning the ego system have been introduced in (1). The constraint (15e) refers to the safety requirement on the distance between the ego system and obstacle $s$, where ${\rm dis}(\cdot)$ is a distance measure between $p^{e}_{i|t}$ the position of center of geometry of the ego system and the obstacle’s occupancy $\hat{\mathcal{O}}_{i|t}^{s}$ [34]. It means at time step $t+i$ in the prediction horizon, the distance between $p^{e}_{i|t}$ and $\hat{\mathcal{O}}_{i|t}^{s}$ should be larger than the minimum safety distance $d^{s}_{\rm min}$. Note that $p_{i|t}^{e}$ and $\hat{\mathcal{O}}_{i|t}^{s}$ predict the motion of the center of geometry of the ego and obstacle systems, such that the actual shape factors of them should be considered in designing $d^{s}_{\rm min}$. In addition, (15e) can be designed with a slack variable to enhance the feasibility. This will be introduced in Example V.1 with the explicit and equivalent expressions of constraint (15e).
The motion planning of the ego system is performed by solving the OCP (V) at every time step $t$ to obtain the optimal control sequence $\{u^{e,\star}_{i-1|t}\}_{i=1}^{N}$, which is applied to steer the model (15b) to generate the reference trajectory for a lower-level controller at the current time step.
Example V.1.
MPC formulation of the reach-avoid problem: Consider the models of EV and SV in Example III.1. Based on the occupancy $\hat{\mathcal{O}}_{i|t}^{s}$ of the SV, the EV model (III.1), and formulation (V), the OCP in this case is designed as
$\mathop{\rm minimize}\limits_{u_{i-1|t}^{e},\ \lambda_{i}^{s},\ \varepsilon_{i}^{s}}\quad$ $\sum_{i=1}^{N}\left(||\delta_{i-1|t}^{e}||_{Q_{1}}^{2}+||\eta_{i-1|t}^{e}||_{Q_{2}}^{2}\right)+$
$||E_{N|t}||_{Q_{3}}^{2}+\sum_{i=1}^{N}\left(||\varepsilon_{i}^{s}||_{Q_{4}}^{2}\right)$ (16a)
${\rm subject\ to}\quad$ $x_{i|t}^{e}=f(x_{i-1|t}^{e},\ u_{i-1|t}^{e}),$ (16b)
$\underline{\mathcal{U}}\leq\left[v_{i|t}^{e}\ a_{i|t}^{e}\ \delta_{i-1|t}^{e}\right]^{\rm T}\leq\overline{\mathcal{U}},$ (16c)
${p_{i|t}^{e}}\in\mathcal{D},$ (16d)
$(H_{i|t}^{s}p_{i|t}^{e}-h_{i|t}^{s})^{\rm T}\lambda_{i}^{s}\geq d^{s}_{\rm min}-\varepsilon_{i}^{s},$ (16e)
$||(H_{i|t}^{s})^{\rm T}\lambda_{i}^{s}||_{2}=1,$ (16f)
[…]
where $Q_{1}$,normal-$\cdots$,$Q_{4}$ are weighting matrices, (16b) is the discrete approximation of the continuous model (III.1) using a fourth-order Runge-Kutta method with the sampling interval $T$, $u_{i-1|t}^{e}=[\delta_{i-1|t}^{e}\ \eta_{i-1|t}^{e}]$ is the control input of model (III.1), and $x_{i-1|t}^{e}$ is the state of model (III.1). Parameters $\underline{\mathcal{U}}$ and $\overline{\mathcal{U}}$ mean the lower and upper bounds on the velocity, acceleration, and front tire angle of the EV. Constraint (16d) limits the position of the center of geometry of the EV, $p_{i|t}^{e}=[p_{x,i|t}^{e}\ p_{y,i|t}^{e}]^{\rm T}$, within the driveable area $\mathcal{D}$. Constraints (16e)–(16g) are an equivalent reformulation of the collision-avoidance constraint (15e) with additional decision variable $\lambda_{i}^{s}$ and slack variable $\varepsilon_{i}^{s}$ [34, Section III-B]. The parameter $n_{h}$ is the number of edges of the set $\hat{\mathcal{O}}^{s}_{i|t}$, and the matrix $H^{s}_{i|t}$ and the vector $h^{s}_{i|t}$ are given by
$\hat{\mathcal{O}}^{s}_{i|t}=\left\{x\in\mathbb{R}^{n_{p}}:H^{s}_{i|t}x\leq h^{s}_{i|t}\right\}.$
Define $l^{e}$ and $w^{e}$ as the length and width of the EV, $l^{s}$ and $w^{s}$ the length and width of the SV. The safety distance $d^{s}_{\rm min}$ in (16e) is designed as
$d^{s}_{\rm min}=\sqrt{(l^{e}/2)^{2}+(w^{e}/2)^{2}}+\sqrt{(l^{s}/2)^{2}+(w^{s}/2)^{2}},$
Given $v^{\rm f}$, $x^{\rm f}$, $y^{\rm f}$, and $\varphi^{\rm f}$ as the predefined reference velocity, longitudinal and lateral positions, and heading angle of the EV, the terminal error vector $E_{N|t}$ in (16a) is defined as
$E_{N|t}=[v^{e}_{N|t}-v^{\rm f}\ x^{e}_{N|t}-x^{\rm f}\ y^{e}_{N|t}-y^{\rm f}\ \varphi^{e}_{N|t}-\varphi^{\rm f}].$ (17)
VI Simulations of Reach-Avoid Motion Planning
[…]
TABLE II: General Parameters in Simulations
Symbol Value Symbol Value
$l_{f}$, $l_{r}$ $0.08\ {\rm m}$, $0.08\ {\rm m}$ $T$ $0.25\ {\rm s}$
$l^{e}$, $w^{e}$ $0.26\ {\rm m}$, $0.25\ {\rm m}$ $l^{s}$, $w^{s}$ $0.36\ {\rm m}$, $0.23\ {\rm m}$
$N$ 10 $Q_{1}$, $Q2$ $1$, $1$
[…]
This section presents simulation results for reach-avoid planning as shown in Fig. 1, where the EV and SV are specified as a car-like mobile robot in this case. The implementations can be found in our published code [^2] https://github.com/JianZhou1212/robust-mpc-motion-planning-by-learning-obstacle-uncertainties. The mission of the EV is to follow a predefined reference state while avoiding collision with the SV and borders of the driveable area. It is assumed that the SV holds a higher priority, requiring the EV to actively take measures to avoid collisions. The controller of the SV is simulated by a nonlinear MPC to track its reference target, where the reference target and the model and controller parameters are unknown to the EV. The SV takes control input from an intended control set $\tilde{\mathbb{U}}_{t}^{s}$ to plan a reference trajectory to reach its reference target; the set $\tilde{\mathbb{U}}_{t}^{s}$ and the distribution of control actions of the SV are also unknown to the EV. From the perspective of the EV, the SV is modeled by (4).
In the numerical studies, the computation of the occupancy of the SV firstly requires formulating an admissible control set $\mathbb{U}^{s}$ according to Assumption IV.2. Next, the LP problem (13) is solved to learn the unknown intended control set of SV. This is initialized by an initial information set $\mathcal{I}_{0}^{s}$ that contains some artificial samples to construct a small but non-empty set $\hat{\mathbb{U}}_{0}^{s}$ by solving (8). This implies that the proposed method does not have any prior information on the uncertainties of the SV. Then, for $t\in\mathbb{N}_{+}$ the set $\hat{\mathbb{U}}_{t}^{s}$ is obtained by solving (13) based on $\hat{\mathbb{U}}_{t-1}^{s}$ and $u_{t-1}^{s}$, where $\hat{\mathbb{U}}_{t-1}^{s}$ is obtained at time step $t-1$, and $u_{t-1}^{s}$ contains the measured longitudinal and lateral accelerations of the SV in the ground coordinate system at time step $t-1$. The learned set $\hat{\mathbb{U}}_{t}^{s}$ is substituted into (IV-C) to obtain $\hat{\mathcal{O}}_{i|t}^{s}$, i.e., the predicted occupancy of the SV, for planning of the EV by solving (V.1).
The proposed planner is compared with a robust MPC (RMPC) and a deterministic MPC (DMPC), where the RMPC and DMPC are implemented by replacing $\hat{\mathbb{U}}_{t}^{s}$ in (14b) with $\mathbb{U}^{s}$ and ${\bm{0}}$, respectively. The set addition and polytope computations were implemented by the Python package pytope [35]. The optimization problems involved were solved by CasADi [36] and Ipopt [37] using the linear solver MA57 [38]. The simulations were performed on a standard laptop running Ubuntu 22.04 LTS and Python 3.10.12. Table II provides a summary of simulation parameters.
Figure 8: Comparison of $d^{e,s}$ and $d^{e,\rm ref}$ during the motion. (a) The polytope distance between EV and SV ($d^{e,s}$) during the motion (the minimum EV-SV distance with DMPC is $0.03\ {\rm m}$, and that with the proposed method is $0.14\ {\rm m}$). (b). The Euclidean distance between EV and its reference target ($d^{e,\rm ref}$) during the motion. (c). The cost function value.
TABLE III: Comparison of the Performance of Three Planners in Stochastic Simulations
Planner & success rate ${\rm min}(d^{e,s})$ $\tau^{e,\rm ref}$ Sum of Cost
Name Collision-free rate Complete rate Mean Min. Mean Max. Mean Max.
Proposed $100\%$ $100\%$ $0.104$ $0.062$ $8.75$ $8.75$ $2000$ $2331$
RMPC $100\%$ $73\%$ $0.711$ $0.199$ $11.26$ $13.75$ $4526$ $6357$
DMPC $76\%$ $100\%$ $0.013$ $5.7{\rm e}^{-4}$ $8.75$ $8.75$ $1966$ $1980$
[…]
The methods are first implemented in a single case to get insights into the performance of each planner. The comparison is based on the planned paths, the polytope distance between the EV and SV ($d^{e,s}$), the Euclidean distance between the EV and its reference state ($d^{e,\rm ref}$), which reflects the convergence speed to the reference, and the cost function value. The results are shown in Fig. 7 and Fig. 8, respectively. It is observed in Fig. 7 that the three methods perform differently in the same scenario as a result of different considerations of uncertainties. Further analysis from Fig. 8(a) shows that the minimum distance between EV and SV with DMPC is close to $0\ {\rm m}$, and that with the proposed method is $0.14\ {\rm m}$, which is sufficient in this case considering that the length of the EV is $0.26\ {\rm m}$. In contrast, RMPC can generate the safest reference trajectory as it maintains a larger distance between EV and SV. However, the cost of pursuing robustness is that the EV needs a longer time and a larger cost to converge to the reference state, as shown in Figs. 8(b)–(c), while the DMPC and the proposed method converges faster with a smaller cost.
The three methods are further compared through $100$ Monte-Carlo simulations with a randomly sampled initial state for the SV, while other conditions are the same as in Fig. 7. The planners are compared by counting the collision-free rate of planning, where a collision-free mission is fulfilled if the EV has no collision with either the SV or the driving-area boundary. Among all successful cases, the completed cases are counted, where a complete mission means that the Euclidean distance between the state of the EV and its reference state converges to less than or equal to $0.2$ within a predefined time limitation ($13.75\ {\rm s}$ in the simulation). For all completed cases with each planner, the minimum polytope distance between the EV and SV of each simulation (${\rm min}(d^{e,s})$), the first time instant when the EV reaches the reference state ($\tau^{e,\rm ref}$), and the sum of the cost function value are compared, and the results are presented in Table III. It is seen that the proposed method and RMPC achieve a collision-free rate of $100\%$, whereas DMPC achieves $76\%$. Among all collision-free cases, the proposed method and DMPC have a $100\%$ complete rate, while the complete rate for RMPC is $73\%$. Analyzing the random variables ${\rm min}(d^{e,s})$, $\tau^{e,\rm ref}$, and the sum of cost reveals that both the proposed method and DMPC guide the EV to the reference state at the same time instant, and the costs are close. However, the proposed method demonstrates superior safety compared to DMPC. On the other hand, RMPC ensures a large distance between the EV and SV in each simulation, but it takes more time and a larger cost to converge.
The average execution time for each iteration of the proposed method, DMPC, and RMPC is $0.105\ {\rm s}$, $0.056\ {\rm s}$, and $0.103\ {\rm s}$, respectively, with standard deviations of $0.025\ {\rm s}$, $0.023\ {\rm s}$, and $0.022\ {\rm s}$, respectively. These results were recorded from $10$ random simulations where each MPC algorithm was run $55$ steps. The simulation results demonstrate that the proposed method based on learning obstacle uncertainties performs successfully in the absence of prior knowledge of the environment and the obstacle. It is as safe as the RMPC, while simultaneously enhancing efficiency and reducing the cost associated with motion-planning tasks. In contrast, when compared to the DMPC, it markedly increases safety without incurring additional costs or compromising the efficiency of executing the tasks.
Remark VI.1.
The DMPC and RMPC are selected as the baseline methods for two primary reasons. First, in the studied reach-avoid problem, the DMPC excels in planning the most efficient reference trajectory when the motion-planning problem is feasible, while the RMPC is known for its guaranteed safety. The evaluations of the proposed method are demonstrated by comparing it with the efficiency-optimal method and the safety-optimal method, respectively. Second, similar to the proposed method, the DMPC and RMPC rely on a few assumptions on the traffic scenario. Other methods such as risk-aware optimal control [4], interaction and safety-aware MPC [24], and non-conservative stochastic MPC [23], [27], require certain assumptions on uncertainty distributions of the SVs, which are not satisfied in our case studies. Therefore, DMPC and RMPC provide a fair comparison for assessing the proposed method’s performance.
Remark VI.2.
In the implementations, one important step is to, according to Assumption IV.1, estimate the control input $u_{t-1}^{s}$ of model (4) of the SV at every time step $t$. This is achieved by observing the velocities of the SV at time steps $t-1$ and $t$ to calculate its acceleration at time step $t-1$ in the vehicle frame, which is combined with its observed orientation at time step $t-1$ to estimate $u_{t-1}^{s}$.
Remark VI.3.
The case studies consider the motion-planning problem for the EV in the presence of one SV, while the method can also manage multiple obstacles. The effectiveness of the proposed method in scenarios involving multiple obstacles has been verified in simulation, and the results are provided together with the videos in Section VII-B.
VII Hardware Experiments
Figure 9: The experiment scenarios of motion planning of an EV moving along a track with an SV.
TABLE IV: Experimental Parameters in Each Scenario
Scenario Initial States Ref. V. of EV Ref. V. of SV
Encounter EV: $2.50\ {\rm m}$, $0\ {\rm m}$, $\pi/2\ {\rm rad}$, $0\ {\rm m/s}$ SV: $-2.50\ {\rm m}$, $0\ {\rm m}$, $\pi/2\ {\rm rad}$, $0\ {\rm m/s}$ $0.45\ {\rm m/s}$ $0.45\ {\rm m/s}$
Overtaking EV: $2.50\ {\rm m}$, $-0.40\ {\rm m}$, $\pi/2\ {\rm rad}$, $0\ {\rm m/s}$ SV: $1.60\ {\rm m}$, $2.50\ {\rm m}$, $\pi\ {\rm rad}$, $0\ {\rm m/s}$ $0.45\ {\rm m/s}$ $0.25\ {\rm m/s}$
Evasion EV: $1.50\ {\rm m}$, $2.50\ {\rm m}$, $\pi\ {\rm rad}$, $0\ {\rm m/s}$ SV: $2.50\ {\rm m}$, $0.90\ {\rm m}$, $\pi/2\ {\rm rad}$, $0\ {\rm m/s}$ $0.35\ {\rm m/s}$ $0.70\ {\rm m/s}$
[…]
Figure 10: Measured trajectories of EV and SV during the encounter scenario. (a) The first encounter (from $8.28\ {\rm s}$ to $15.42\ {\rm s}$). (b) The second encounter (from $26.33\ {\rm s}$ to $33.92\ {\rm s}$).
Figure 11: Information about SV during the encounter scenario. (a) The learned set $\hat{\mathbb{U}}^{s}_{t}$ over time in the encounter experiment (the subfigure indicates the final set $\hat{\mathbb{U}}^{s}_{t}$). (b)The velocity of SV in the ground coordinate system.
Figure 12: Visualization of the predicted occupancy $\hat{\mathcal{O}}_{i|t}^{s}$ at the initial moments in the encounter scenario.
Figure 13: Visualization of the predicted occupancy of the SV and the planned path of the EV at key moments during the encounter (left column), and the corresponding scenarios in the experiment (right column).
Figure 14: Measured trajectories of EV and SV during the overtaking experiment (from $9.56\ {\rm s}$ to $17.50\ {\rm s}$).
Figure 15: Measured trajectories of EV and SV during the evasion experiment (from $6.65\ {\rm s}$ to $11.27\ {\rm s}$).
This section presents hardware-experiment results in scenarios different from those in Section VI to demonstrate the method’s applicability across different situations, as well as the real-time performance of the method in applications. The experiment videos are accessible online [^3] https://youtu.be/r8YCDAh7Hts.
VII-A Scenario Description
The experimental environment is illustrated in Fig. 9, where the EV and SV are represented by a hamster platform and a cascar platform, respectively. The size parameters of the EV and SV are the same as those in Table II. Both platforms are kinematically equivalent to a full-scale vehicle and have similar dynamic characteristics, including suspension and tire systems, as those of a regular vehicle. This makes them suitable platforms for testing the proposed method in realistic scenarios. In the experiments, the mission of the SV is to move along the center line of the track with a predefined reference velocity, and the objective of the EV is to track the center line of the track and a predefined reference velocity while ensuring collision avoidance with the SV. The SV holds a higher priority than the EV, such that the EV needs to apply the proposed method to replan the reference trajectory at every time step. The SV takes control inputs from an unknown control set $\tilde{\mathbb{U}}_{t}^{s}$. The motion-planning performance of the EV is evaluated in three different scenarios:
  • Encounter: EV and SV move in opposite directions, with the EV avoiding collisions with the SV.
  • Overtaking: the faster rear EV surpasses a slower preceding SV.
  • Evasion: the slower preceding EV takes evasive action to avoid a collision with a faster rear SV.
The design of the MPC planner follows the same approach as presented in Section VI, so details are omitted for conciseness. Note that to enhance the efficiency of solving the OCP problem, several adaptations are applied. First, the dimension of the EV model in (III.1) is reduced to $4$ by taking the longitudinal acceleration and front tire angle as the inputs. Second, the prediction horizon is changed to $N=9$.
VII-B Experiment Results and Discussion
In the experiments, the polytope computation and the optimization problem are solved by the same tools as in Section VI. The positioning of both EV and SV was performed by the Qualisys motion-capture system [39], and the control inputs of the vehicles were computed on a laptop running Ubuntu 20.04 LTS operating system. The communications were established by the Robot Operating System (ROS Noetic). The reference paths of the EV and SV were followed by a pure-pursuit controller [40], and the reference velocities were followed by the embedded velocity-tracking controllers of the hamster platform and the cascar platform, respectively. The initial states and reference velocities of the EV and SV in each scenario are collected in Table IV. The reference velocity for the hamster platform is designed following the specifications detailed in [41] and [42].
In the encounter scenario, the EV and SV will intersect twice. The trajectories of both vehicles during these encounters are depicted in Fig. 10. Fig. 11(a) visualizes the learned control set $\hat{\mathbb{U}}_{t}^{s}$ of the SV over time, and Fig. 11(b) presents the velocity of the SV in the ground coordinate system. Following Fig. 11(a), Fig. 12 illustrates how the predicted occupancy of the SV over the prediction horizon, i.e., $\hat{\mathcal{O}}_{i|t}^{s}$, evolves at the initial moments. Leveraging the occupancy $\hat{\mathcal{O}}_{i|t}^{s}$, the EV plans its reference by solving problem (V). The occupancy $\hat{\mathcal{O}}_{i|t}^{s}$, along with the planned reference path of the EV at key moments during the first encounter, are exhibited in Fig. 13.
In the encounter experiments illustrated in Fig. 10, the EV successfully avoids a collision with the SV. Fig. 11(a) demonstrates that the set $\hat{\mathbb{U}}^{s}_{t}$ expands over time as more information is obtained, ultimately converging after approximately $4\ {\rm s}$. This observation aligns with the SV’s velocity, as seen in Fig. 11(b), which stabilizes after $4\ {\rm s}$. Fig. 12 further illustrates the impact of the set $\hat{\mathbb{U}}^{s}_{t}$ on the predicted occupancy $\hat{\mathcal{O}}_{i|t}^{s}$, which increases at the initial moments and subsequently tends to stabilize when $\hat{\mathbb{U}}^{s}_{t}$ ceases to expand. This reflects that in the online execution, the robustness of the method strengthens as the information set $\mathcal{I}_{t}^{s}$ becomes richer. Finally, the results presented in Fig. 13 affirm that the EV can effectively plan collision-avoidance reference trajectories in the presence of the uncertain dynamic obstacle, and move back to the track when the obstacle no longer has the risk of collision.
The trajectories of the EV in the overtaking and evasion experiments are illustrated in Fig.14 and Fig.15, respectively. In both scenarios, the proposed method successfully addresses the motion-planning challenges. The corresponding learned sets $\hat{\mathbb{U}}_{t}^{s}$ and the predicted occupancies $\hat{\mathcal{O}}_{i|t}^{s}$, along with the planned collision-avoidance trajectories of the EV, exhibit qualitative similarities. In the presented experiments, the average execution time of the proposed method is $0.064\ {\rm s}$, and the standard deviation is $0.014\ {\rm s}$, with the motion planner executed $200$ times in the encounter scenario. The computation time in the overtaking and evasion scenarios is comparably similar. Both the simulations and experiments demonstrate the method’s effectiveness in addressing environmental uncertainties and provide validation its real-time applicability.
VIII Conclusion and Future Work
This paper studied safe motion planning of autonomous robotic systems under obstacle uncertainties. An online learning-based method was proposed to learn the uncertain control set of the obstacles to compute the forward reachable sets, which were integrated into a robust MPC planner to obtain the optimal reference trajectory for the ego system. Simulations and hardware experiments with an application of the method to an autonomous driving system in different scenarios show that: (1) The method is safer than deterministic MPC and less conservative than worst-case robust MPC, while maintaining safety in uncertain traffic environments; (2) The method can perform safe motion planning in uncertain environments without prior knowledge of uncertainty information of the obstacles; (3) The method is applicable and real-time implementable in practical scenarios.
Future research will focus on extending the proposed method for nonlinear obstacle models and other environmental uncertainties. Another research direction of great interest is the interaction-aware motion planning of multi-agent systems by proactively learning the decision-making of other agents.
Acknowledgments
The authors would like to acknowledge Prof. Lars Nielsen at Linköping University for discussions regarding the second-order integrator and polytope-based collision-avoidance constraints. Lic. Theodor Westny at Linköping University is acknowledged for skillful assistance in the experiments.